I know the base of my triangle is 3.8m and the height is 1.4m. Can I work out how long the sides are(they are equal length). I dont know any of the angles.
Answer by stanbon(75887)

I know the base of my triangle is 3.8m and the height is 1.4m. Can I work out how long the sides are(they are equal length). I dont know any of the angles.
------------------
The height is perpendicular to the base so you have two right
triangles.
The height bisects the base so you have 2 base segment of 1.9m each
------------------
Use Pythagoras to find the length of a side:
s^2 = 1.9^2 + 1.4^2
s^2 = 5.57
side = 2.36
=====================
Cheers,
Stan H.